# Turning English-centric LLMs Into Polyglots: How Much Multilinguality Is Needed?

### v1

@@ Abstract
The vast majority of today’s large language models are English-centric, having been pretrained predominantly on English text. Yet, in order to meet user expectations, models need to be able to respond appropriately in multiple languages once deployed in downstream applications. Given limited exposure to other languages during pretraining, cross-lingual transfer is important for achieving decent performance in non-English settings. In this work, we investigate just how much multilinguality is required during finetuning to elicit strong cross-lingual generalisation across a range of tasks and target languages. We find that, compared to English-only finetuning, multilingual instruction tuning with as few as three languages significantly improves a model’s cross-lingual transfer abilities on generative tasks that assume input/output language agreement, while being of less importance for highly structured tasks. Our code and data is available at https://github.com/ZurichNLP/multilingual-instruction-tuning.
@@
[…]
@@ 1 Introduction
Conversational instruction tuning is a popular method for aligning large pretrained language models (LLMs) with user expectations such that they can effectively respond to a user’s input query and follow natural language instructions Ouyang et al. (2022); Wang et al. (2023); Chiang et al. (2023). An implicit expectation of conversational chatbots is that the language of a model’s response should match that of the user’s input query. For instance, unless otherwise specified, a German-language input should result in a German-language output. However, since the vast majority of pretraining and tuning data used to develop these models is in English, many instruction-tuned LLMs struggle to respond consistently in other languages Touvron et al. (2023b); Chen et al. (2023); Ye et al. (2023); Zhang et al. (2023b).
@@
Figure 1: Input/output (IO) language agreement for English (en), German (de), Bulgarian (bg) and Icelandic (is) when instruction tuning on monolingual English (Mono) or on multilingual data (Multi-Guanaco). Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
@@
Despite this limited exposure, English-centric LLMs such as Llama 2 can achieve near-perfect input/output (IO) language agreement when tuned with relatively few multilingual conversational instructions. Figure 1 depicts IO language agreement, as measured by OpenLID [^1] https://github.com/laurieburchell/open-lid-dataset Burchell et al. (2023) and compares multilingual tuning on the language-diverse Guanaco dataset Dettmers et al. (2023); Köpf et al. (2023) with monolingual tuning on an English-only subset of instructions. As can be seen, multilingual tuning elicits strong IO language agreement, well within the bounds of language identification error rates, for non-English languages seen to differing degrees during pretraining and finetuning without degrading performance on English.
@@
This observation raises two major questions, which we aim to address in this paper: Q1: How much multilinguality is required during finetuning to elicit few- and even zero-shot cross-lingual generalisation in English-centric LLMs? Q2: Which languages and tasks benefit most from multilingual instruction tuning of English-centric LLMs?
@@
To investigate these questions, we instruction-tune English-centric LLMs, varying the number of languages seen during finetuning and evaluate performance in multiple target languages and on different tasks. We consider both high- and low-resource language settings with different scripts and investigate performance on generative tasks, such as open-ended chat and extractive question answering, as well as structured tasks aimed at assessing commonsense reasoning and language understanding.
@@
Our results indicate that multilingual instruction tuning is crucial for eliciting cross-lingual transfer on generative tasks that assume IO language agreement while being less important in structured tasks that are commonly used to benchmark LLM performance. Furthermore, we empirically show that only a small number of finetuning languages is required to promote cross-lingual transfer. This highlights that tuning data for all potential target languages is not necessary to derive a capable polyglot chat model from an English-centric LLM.
@@ 2.1 Instruction Tuning LLMs
Instruction tuning describes a supervised finetuning (SFT) strategy that aims to provide a model with a diverse set of demonstrations of user input queries paired with desirable model outputs. Unlike task-specific SFT, instruction tuning aims to promote cross-task generalisation, allowing for a ‘generalist’ model that is capable of completing any text-based task on the basis of natural language instructions provided at inference time Mishra et al. (2022); Wei et al. (2022); Wang et al. (2022); Sanh et al. (2022); Longpre et al. (2023). Meanwhile, framing instructions in a conversational manner and over multiple dialogue turns has been shown to be effective at deriving performant chat models Taori et al. (2023); Conover et al. (2023); Chiang et al. (2023); Dettmers et al. (2023); Ding et al. (2023). Furthermore, LLM instruction tuning remains effective given relatively limited labelled data (Ouyang et al., 2022; Touvron et al., 2023b; Zhou et al., 2023), parameter efficient methods Hu et al. (2021); Zhang et al. (2023a) and model quantisation Dettmers et al. (2023); Li et al. (2023).
@@ 2.2 Cross-lingual Transfer in English-centric LLMs
Despite numerous pushes towards improving NLP support for non-English and low-resource languages Costa-jussà et al. (2022); Le Scao et al. (2023), the vast majority of today’s publicly available LLMs are English-centric. For instance, GPT-3’s training data consisted of approximately 93% English documents with the remaining 7% pertaining to other languages Brown et al. (2020). [^2] https://github.com/openai/gpt-3/blob/master/dataset_statistics This trend is further reflected in popular open-weight LLMs (see Table 1). One potential reason for this could be the “the curse of multilinguality” Conneau et al. (2020) which describes how as multilinguality increases, evaluation performance decreases. For example, multilingual LLMs have been shown to sacrifice performance on English tasks due to having to share a finite amount of pre-training tokens and model parameters across more languages Lin et al. (2022); Le Scao et al. (2022). Consequently, the cross-lingual transfer abilities of more performant, English-centric models is important for many NLP practitioners looking to deploy these models in real-world settings.
A number of works have reported on the multilingual abilities of proprietary models such as GPT-3 and its derivatives across a range of NLU and NLG benchmarks Lai et al. (2023a); Holmström et al. (2023); Armengol-Estapé et al. (2022). On tasks such as machine translation (MT), GPT-3.5-Turbo has been shown to outperform dedicated systems for some high-resource languages Hendy et al. (2023); Lu et al. (2023); Jiao et al. (2023); Bang et al. (2023), particularly when translating between two non-English languages Laskar et al. (2023). However, performance on low-resource languages generally lags behind.
Meanwhile, other studies have focused on open-weight LLMs. Ye et al. (2023) compare multilingual reasoning capabilities of pretrained BLOOM Le Scao et al. (2023) and LLaMA Touvron et al. (2023a) and find that despite minimal amounts of non-English pretraining data, LLaMA has strong cross-lingual transfer abilities. Muennighoff et al. (2023) find that multilingual multitask finetuning improves BLOOM’s zero-shot abilities across all target languages compared with English-only finetuning. Finally, Chen et al. (2023) show that large-scale multilingual instruction tuning can improve performance on open-ended chat in multiple target languages.
This last work is most closely related to our current work but differs in a few key factors. Firstly, we investigate multilingual chat capabilities of more recent English-centric models, pretrained with significantly more data. Secondly, inspired by Zhou et al. (2023) we adopt a less-is-more finetuning approach, controlling the amount and distribution of multilingual data. In doing so, we focus our analysis on the minimal amount of multilinguality needed to elicit cross-lingual transfer. Finally, our evaluations consider more target languages and distinct downstream tasks.
@@ 3 Experimental Setup
To explore the multilingual capabilities of English-centric LLMs, we instruction-tune a series of models on a fixed-size set of examples, varying the number of languages available. Following this, we evaluate the resulting models in multiple target languages on four distinct tasks that are representative of how LLMs may be used in downstream applications.
@@ 3.1 English-centric LLMs
A prevailing trend in the development of recent LLMs is a clear focus on scaling up the size of the pretraining corpus Hoffmann et al. (2022). For instance, open-weight LLMs such as Falcon Almazrouei et al. (2023) and Llama 2 Touvron et al. (2023b) were trained on 1.5 and 2 trillion tokens respectively. Yet, although these numbers far surpass the 300 billion tokens used to train GPT-3 Brown et al. (2020), the distribution of language coverage remains similar across models with more than 90% pertaining to English (see Table 1). For our main experiments, we focus on Llama 2 7b, but also consider Llama 2 70b (§5.2) and Falcon 7b (Appendix E) to study the effect of model scaling and different training approaches respectively.
@@ 3.2 Instruction-tuning Data
For instruction tuning, we take inspiration from Dettmers et al. (2023) and finetune on high-quality conversations from the OpenAssistant dataset Köpf et al. (2023). These conversations comprise multiple dialogue turns between crowdworkers who were asked to either interact with or assume the role of a helpful AI assistant.
In contrast to Dettmers et al. (2023), who use all 9,846 top-rated conversations to train their ‘Guanaco’ models, we subsample training instances from the Guanaco dataset in order to control the amount of multilinguality. Specifically, we sample 3,200 unique English instances as an initial monolingual dataset, which we refer to as ‘Mono’. To construct datasets for multilingual finetuning we sample 200 unique training examples from each of five most frequent non-English languages in Guanaco (Spanish, Russian, German, Chinese, French). Given these subsets, we incrementally substitute English examples in Mono for non-English ones, one language at a time, following the order of how frequently each language appears in Guanaco. The resulting multilingual datasets are denoted as Multi-$i$, where $i$ equals the number of distinct languages included. For comparison, we also train models on the full Guanaco dataset, which includes more than 30 distinct languages.
@@ 3.3 Instruction-tuning and Inference Settings
For efficient instruction tuning, we use LoRA adapters ($R=64$, $\alpha=16$), leveraging Hugging Face’s TRL library [^3] https://github.com/huggingface/trl. We train all models for 2,000 update steps, using an effective batch size of 64 and a learning rate of $1^{e-5}$. Sequences longer than 1024 tokens are truncated. For 7-billion parameter models, we use two NVIDIA GeForce RTX 3090 with 24GB of memory. The time required for each training run is approximately 8 hours. For the larger 70-billion parameter model, discussed in §5.2, we use the same hyperparameters and train on four NVIDIA A100 GPUs with 80GB of memory each. Here, a single training run takes approximately 20 hours.
At inference time, we load the instruction-tuned models with vLLM Kwon et al. (2023). For open-ended generation tasks, we use top-p sampling (p=0.9) with a temperature of 0.8 and frequency penalty of 0.2. For the more constrained generation tasks we reduce the temperature to 0.001. To account for randomness under these sampling parameters, we report aggregated results for all tasks by repeating inference experiments with three random seeds.
@@ 3.4 Evaluation Tasks and Languages
As evaluation tasks, we consider open-ended chat, extractive question answering, commonsense reasoning and natural language inference. As all of these tasks differ in terms of the availability and representation of ground-truth labels, we describe the specific evaluation strategies in the following section.
Finally, we investigate performance on different target languages, which we select based on the makeup of Llama 2’s pretraining data (see Table 1). In doing so, we aim to investigate i) the supervised setting, where a target language has been seen during pretraining and finetuning; ii) the zero-shot cross-lingual setting, where a target language may have been seen during pretraining but not during finetuning Wu and Dredze (2019); and iii) the extreme low-resource setting, where the target language has not intentionally been seen at any stage of the training procedure. According to Touvron et al. (2023b), German (de), French (fr), Swedish (sv), Chinese (zh), Spanish (es) and Russian (ru) represent the most frequently seen non-English pretraining languages with an estimated 3.4 to 2.6 billion tokens. Catalan (ca), Norwegian (no) and Bulgarian (bg) represent languages less frequently seen during pretraining with an estimated 800 to 400 million tokens. And finally, we select Icelandic (is), Hindi (hi) and Greek (el) to represent extremely low-resource languages, whose frequency in Llama 2’s pretraining data is not available but may still appear in small amounts due to intentional or unintentional contamination Blevins and Zettlemoyer (2022). Unfortunately, existing benchmark evaluation datasets do not cover all of these languages and thus we limit the target languages in those tasks to the available subset of our target languages.
@@ 4 Experiments
4.1 Chat
General-purpose chatbots are a popular application of instruction-tuned LLMs. To evaluate chat performance, we make use of the AlpacaEval prompt dataset Dubois et al. (2023), which includes a diverse set of prompts for open-ended questions, creative writing, brainstorming, and other tasks. We randomly sample 300 prompts and translate these into each target language. As a translation engine, we follow Lai et al. (2023b) and use GPT-3.5-Turbo [^4] To translate chat prompts from AlpacaEval, we use gpt-3.5-turbo-0613.. In contrast to using dedicated translation systems, employing GPT-3.5-Turbo for this purpose has the advantage of being able to explicitly specify instructions that allow for preserving code blocks, tables and terminology, which we include as part of our translation prompt (see Figure B). Furthermore, since GPT-3.5-Turbo is trained on instruction- and conversational-style data, we expect it to perform well at translating in this domain.
To automatically evaluate open-ended chat responses, we leverage LLM-as-a-judge (Zheng et al., 2023). Following Zhou et al. (2023), given an input prompt and model’s response, we ask GPT-3.5-Turbo [^5] For chat evaluation, we use gpt-3.5-turbo-1106 due to its longer context window and cheaper inference costs. to grade the helpfulness of the response on a 6-point Likert scale (see Figure 16 for the prompt used). For each evaluation instance, we provide the prompt and model-generated response directly in the target language, which we found to be on par with evaluating via first translating responses into English (cf. Hada et al., 2023) (see Appendix C for more details). As noted by Chen et al. (2023), GPT-3.5-Turbo sometimes ignores the fact that the output language differs from the input language. Therefore, we force a score of 1 (indicating least helpful) if the language of the response does not match the intended target language.

### v2

@@ Abstract
The vast majority of today’s large language models (LLMs) are English-centric, having been pretrained predominantly on English text. Yet, in order to meet user expectations, models need to be able to respond appropriately in multiple languages once deployed in downstream applications. This requires strong cross-lingual transfer abilities. In this work, we investigate the minimal amount of multilinguality required during finetuning to elicit cross-lingual generalisation in English-centric LLMs. In experiments across four LLMs, we find that multilingual instruction tuning with as few as two to three languages is both necessary and sufficient to elicit effective cross-lingual generalisation, with the limiting factor being the degree to which a target language is seen during pretraining. Evaluations on five different tasks further reveal that multilingual instruction tuning is most beneficial for generative tasks that assume input/output language agreement, such as in chat settings, while being of less importance for highly structured classification-style tasks. [^1] Our code and data is available at https://github.com/ZurichNLP/multilingual-instruction-tuning.
@@
Tannon Kew1thanks:   Work partially conducted during a research internship at Textshuttle.        Florian Schottmann2,3       Rico Sennrich1,4 1University of Zurich, 2Textshuttle, 3ETH Zurich, 4University of Edinburgh {kew,sennrich}@cl.uzh.ch, schottmann@textshuttle.com
@@ 1 Introduction
Conversational instruction tuning is a popular method for aligning large pretrained language models (LLMs) with user expectations such that they can effectively respond to a user’s input query and follow natural language instructions Ouyang et al. (2022); Wang et al. (2023); Chiang et al. (2023); Zhou et al. (2023). An implicit expectation of conversational chatbots is that the language of a model’s response should match that of the user’s input query. For instance, unless otherwise specified, a German-language input should result in a German-language output. However, since the vast majority of training data is in English, many instruction-tuned LLMs struggle to respond consistently in other languages Ouyang et al. (2022); Touvron et al. (2023); Chen et al. (2024); Ye et al. (2023); Zhang et al. (2024b).
@@
Figure 1: Input/output (IO) language agreement for English (en), German (de), Bulgarian (bg), and Icelandic (is) given English-only instruction tuningono) or andmultilingual instruction tuning (Multi-Guanaco). Striped bars indicate that the target language is not seen during finetuning (i.e. the 0-shot setting). Error bars show a confidence interval of 95%.
@@
Despite limited exposure, however, English-centric LLMs such as Llama 2 seemingly achieve near-perfect input/output (IO) language agreement when tuned with relatively few multilingual conversational instructions. Figure 1 depicts IO language agreement, as measured by OpenLID Burchell et al. (2023), and compares multilingual tuning on the language-diverse Guanaco dataset Dettmers et al. (2023); Köpf et al. (2023) with monolingual tuning on an English-only subset of instructions. As can be seen, multilingual tuning elicits strong IO language agreement – within the bounds of OpenLID’s error rates – for non-English languages seen to varying degrees during pretraining and finetuning without degrading performance on English.
@@
This observation raises two major questions, which we aim to address in this paper: Q1: How much multilinguality is required during finetuning to elicit 0-shot cross-lingual generalisation in English-centric LLMs? Q2: Which languages and tasks benefit most from multilingual instruction tuning of English-centric LLMs?
@@
To investigate these questions, we instruction tune four distinct English-centric LLMs with varying degrees of multilinguality and evaluate performance on a diverse set of up to 19 languages across five different tasks. Specifically, we consider high-, medium-, and low-resource target languages with regard to a model’s pretraining data distribution and generative tasks, such as single-turn dialogue, sentence simplification, and extractive question answering, as well as more structured tasks aimed at assessing commonsense reasoning and language understanding.
@@
Our results indicate that multilingual instruction tuning is crucial for eliciting cross-lingual transfer on generative tasks that assume IO language agreement while being less important in structured tasks that are commonly used to benchmark LLM performance. Furthermore, we empirically show that only a small number of finetuning languages is required to promote cross-lingual transfer. This highlights that tuning data for all potential target languages is not necessary to derive capable polyglot chat models from English-centric LLMs.
@@ 2.1 Instruction Tuning LLMs
Unlike task-specific finetuning, instruction tuning aims to promote cross-task generalisation, allowing for a ‘generalist’ model that is capable of completing any text-based task on the basis of natural language instructions provided at inference time Mishra et al. (2022); Wei et al. (2022); Wang et al. (2022); Sanh et al. (2022); Longpre et al. (2023). Meanwhile, framing instructions in a conversational manner and over multiple dialogue turns has been shown to be effective at deriving performant chat models Taori et al. (2023); Conover et al. (2023); Chiang et al. (2023); Dettmers et al. (2023); Ding et al. (2023). Furthermore, LLM instruction tuning remains effective given relatively limited labelled data Ouyang et al. (2022); Touvron et al. (2023); Zhou et al. (2023), parameter efficient training strategies Hu et al. (2021); Zhang et al. (2024a) and model quantisation Dettmers et al. (2023); Li et al. (2023b).
@@ 2.2 Cross-lingual Transfer in English-centric LLMs
The vast majority of today’s publicly available LLMs are English-centric. For instance, GPT-3’s training data consisted of approximately 93% English documents with the remaining 7% pertaining to other languages Brown et al. (2020). [^2] https://github.com/openai/gpt-3/blob/master/dataset_statistics This trend is further reflected in popular open-weight LLMs (see Table 1). One potential reason for this could be the “the curse of multilinguality” Conneau et al. (2020) which arises from having to share a finite model capacity across more languages Lin et al. (2022); Le Scao et al. (2022). Consequently, cross-lingual transfer abilities of performant English-centric models is highly valuable.
Despite this, proprietary models such as GPT-3 and its derivatives have shown impressive performance in multilingual settings across a range of NLU and NLG benchmarks Lai et al. (2023a); Holmström et al. (2023); Armengol-Estapé et al. (2022); Hendy et al. (2023); Lu et al. (2023); Jiao et al. (2023); Bang et al. (2023); Laskar et al. (2023). Meanwhile, there is a growing body of research dedicated to studying similar capabilities of open-weight LLMs Bawden and Yvon (2023) and the benefits of multilingual tuning Ye et al. (2023); Muennighoff et al. (2023). For instance, large-scale multilingual instruction tuning has been shown to improve performance on open-ended chat tasks in multiple target languages for both English-centric and multilingual LLMs Chen et al. (2024); Li et al. (2023a); Weber et al. (2024). In concurrent work to ours, Shaham et al. (2024) report that this holds given minimal multilingual instruction tuning with the multilingual PaLM 2 model Anil et al. (2023).
Our work closely relates to that of Chen et al. (2024) and Shaham et al. (2024) insofar as we explore how multilingual instruction tuning can improve cross-lingual transfer in LLMs. However, in contrast to these works, we focus on the minimal amount of multilinguality needed to elicit cross-lingual transfer in open-weight English-centric LLMs, and we evaluate on a range of tasks and target languages that cover the full distribution of a model’s pretraining data.
@@ 3 Experimental Setup
To explore the multilingual capabilities of English-centric LLMs, we instruction tune a series of models on a fixed-size set of examples, varying the number of languages available. Following this, we evaluate the resulting models in multiple target languages on five distinct tasks that are representative of how LLMs may be used in downstream applications.
@@ 3.1 English-centric LLMs
A prevailing trend in the development of recent LLMs is a clear focus on scaling up the size of the pretraining corpus Hoffmann et al. (2022); Sardana and Frankle (2023). For instance, open-weight LLMs such as Falcon Almazrouei et al. (2023) and Llama 2 Touvron et al. (2023) were pretrained on 1.5 and 2 trillion tokens, respectively. Yet, while this far surpasses the 300 billion tokens used to train GPT-3 Brown et al. (2020), the distribution of language coverage remains similar with more than 90% pertaining to English (see Table 1). For our main experiments, we focus on Llama 2 7b. We then consider Llama 2 70b (§5.2) to study the effect of model scaling on multilinguality. Additionally, we also test whether our findings generalise to other LLMs using Falcon 7b and Llama 3 8b MetaAI (2024), which employ different pretraining approaches (see Appendices F and G).
@@ 3.2 Instruction-tuning Data
For instruction tuning, we take inspiration from Dettmers et al. (2023) and finetune on high-quality conversations from the OpenAssistant dataset Köpf et al. (2023). These conversations comprise multiple dialogue turns between crowdworkers who were asked to either interact with or assume the role of a helpful AI assistant. In contrast to Dettmers et al. (2023), who use all 9,846 top-rated conversations to train their ‘Guanaco’ models, we subsample training instances from the Guanaco dataset in order to control the amount of multilinguality. Specifically, we sample 3,200 unique English instances as an initial monolingual dataset, which we refer to as ‘Mono’. To construct datasets for multilingual finetuning we sample 200 unique training examples from each of the five most frequent non-English languages in Guanaco (Spanish, Russian, German, Chinese, and French). Given these subsets, we incrementally substitute English examples in Mono for non-English ones, one language at a time, following the order of how frequently each language appears in Guanaco. The resulting multilingual datasets are denoted as Multi-$i$, where $i$ indicates the number of distinct languages included. Table 3 provides a detailed summary of the makeup of these datasets. For comparison, we also train models on the full Guanaco dataset, which includes more than 30 distinct languages.
@@ 3.3 Instruction-tuning and Inference Settings
[…]
For efficient inference, we employ vLLM Kwon et al. (2023). We use nucleus sampling Holtzman et al. (2019) ($p=0.9$) with a temperature of 0.8 for open-ended generation tasks and a temperature of 0.001 for the more constrained QA-style tasks to encourage more deterministic behaviour. For all tasks, we report the mean performance across three inference runs using different random seeds.
@@ 3.4 Evaluation Tasks and Languages
As evaluation tasks, we consider single-turn dialogue, sentence simplification, extractive question answering, commonsense reasoning and natural language inference. Since all of these tasks differ in terms of the availability and representation of ground-truth labels, we describe the specific evaluation strategies in the following section.
Figure 2: Average helpfulness of single-turn dialogue responses from Llama 2 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
We select target languages based on the makeup of Llama 2’s pretraining data (see Table 1). In doing so, we aim to study the effect of multilingual instruction tuning in both supervised and 0-shot cross-lingual settings Wu and Dredze (2019), and report results on high-, medium-, and low-resource languages with regard to the pretrained LLM, covering distinct language families and scripts. German (de), French (fr), Swedish (sv), Chinese (zh), [^5] Note, Touvron et al. (2023) do not distinguish between Simplified and Traditional Chinese in their reporting. For our purposes, we explicitly use Simplified Chinese. Spanish (es), Russian (ru), and Italian (it) represent high-resource languages with an estimated 3.4 to 2.2 billion tokens. Portuguese (pt), Vietnamese (vi), Korean (ko), Finnish (fi), Hungarian (hu), Norwegian (no), [^6] Again, lacking specification between Norwegian Bokmål and Nynorsk in Touvron et al. (2023), we explicitly consider the former. Bulgarian (bg), and Slovenian (sl) represent medium-resource languages, with an estimated 1.8 to 0.2 billion tokens. And finally, we select Greek (el), Hindi (hi), and Icelandic (is) as low-resource languages, whose frequency in Llama 2’s pretraining data is not known but are likely appear in small amounts due to contamination Blevins and Zettlemoyer (2022). Unfortunately, existing benchmark evaluation datasets do not cover all of these languages and thus we limit the target languages in those tasks to the available subset of our target languages.
@@ 4 Experiments
4.1 Single-turn Dialogue
General-purpose chatbots are a popular application of instruction-tuned LLMs. To evaluate performance in this type of setting, we make use of the AlpacaEval prompt dataset Dubois et al. (2023), which includes a diverse set of prompts for open-ended questions, creative writing, brainstorming, and other tasks. We randomly sample 300 prompts and translate these into each target language. As a translation engine, we follow Lai et al. (2023b) and use GPT-3.5-Turbo. [^7] To translate prompts from AlpacaEval, we use gpt-3.5-turbo-0613. In contrast to using dedicated translation systems, employing GPT-3.5-Turbo for this purpose has the advantage of being able to explicitly specify instructions that allow for preserving code blocks, tables, and terminology, which we include as part of our translation prompt (see Figure B.1). Furthermore, since GPT-3.5-Turbo is trained on instruction- and conversational-style data, we expect it to perform well at translating in this domain.
To automatically evaluate open-ended responses, we leverage LLM-as-a-judge Zheng et al. (2023). Following Zhou et al. (2023), given an input prompt and model’s response, we ask GPT-3.5-Turbo [^8] For evaluation, we use gpt-3.5-turbo-1106 due to its longer context window. to grade the helpfulness of the response on a 6-point Likert scale (see Figure 20 for the prompt used). For each evaluation instance, we provide the prompt and model-generated response directly in the target language, which we found to be on par with evaluating via first translating responses into English (cf. Hada et al., 2024, see Appendix E for more details). As noted by Chen et al. (2024), GPT-3.5-Turbo sometimes ignores the fact that the output language differs from the input language. Therefore, we force a score of 1 (indicating least helpful) if the language of the response does not match the intended target language according to OpenLID Burchell et al. (2023).